# Implications of recent Higgs data for the light Higgs bosons in the GNMSSM

### v2

@@ 2.3 Trilinear Higgs couplings in the GNMSSM
For the light Higgs scenario studied in this paper, we assume a light CP-odd Higgs boson ($a_{1}$) satisfying $2m_{a_{1}}<m_{h}$, enabling the exotic decay $h\to a_{1}a_{1}$ (i.e. $h\to A_{s}A_{s}$) in both $h_{1}$ and $h_{2}$ scenarios. For the $h_{2}$ scenario, where the lightest CP-even Higgs boson $h_{1}$ is singlet-dominant, $h\to h_{1}h_{1}$ (i.e. $h\to h_{s}h_{s}$) opens when $2m_{h_{1}}<m_{h}$. Both ATLAS and CMS collaborations have performed comprehensive searches for non-SM Higgs bosons $H$, $h_{s}$, $A_{H}$ and $A_{s}$, establishing stringent exclusion bounds on key parameters including masses and relative couplings [54, 55].
The branching ratios of exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ are governed by their respective trilinear Higgs couplings $C_{ha_{1}a_{1}}$($\equiv C_{hA_{s}A_{s}}$) and $C_{hh_{1}h_{1}}$ ($\equiv C_{hh_{s}h_{s}}$). In the GNMSSM, these couplings scale proportionally to vacuum expectation values ($v_{u}$, $v_{d}$, $v_{s}$) or soft SUSY-breaking trilinear terms ($A_{\lambda}$, $A_{\kappa}$). While Ref. [18] provides general expressions, we explicitly define the specific trilinear couplings relevant to our analysis below [28]:
@@
where the last terms in the preceding equations, suppressed by Higgs mixings, depend on $\lambda,\kappa,\tan\beta,v_{s},A_{\lambda}$ and $m_{N}$.
@@ 3 Numerical Result
This section details our scanning strategy and presents numerical results. We construct the GNMSSM model file using SARAH [56, 57, 58, 59] and generate particle spectra with SPheno [60, 61, 62, 56]. For parameter space exploration, we employ the parallelized MultiNest algorithm [63], which utilizes nested sampling to efficiently identify high-likelihood regions, detect multiple solution modes, and compute Bayesian evidence.
@@
$\kappa$ Flat 0.750.75-0.75- 0.750.750.750.750.75 $\tan\beta$ Flat 555560606060
$\lambda$ Flat 000.750.750.750.75 $v_{s}/{\rm TeV}$ Flat 0.10.10.10.11.01.01.01.0
$\delta$ Flat 1.01.0-1.0- 1.01.01.01.01.0 $m_{N}/{\rm TeV}$ Flat 1.01.0-1.0- 1.01.01.01.01.0
$m_{B}/{\rm GeV}$ Flat 1111300300300300 $m_{C}/{\rm GeV}$ Flat 1.01.01.01.0300300300300
$A_{t}/{\rm TeV}$ Flat 1.01.01.01.03.03.03.03.0 $\mu_{\rm tot}/{\rm TeV}$ Flat 0.20.20.20.21.01.01.01.0
$A_{\kappa}/{\rm TeV}$ Flat 2.02.0-2.0- 2.02.02.02.02.0
Table 1: Parameter space explored in this study. All input parameters adopt flat distributions based on their unambiguous physical interpretations. Given the substantial radiative corrections induced by the third-generation squark trilinear couplings ($A_{t}$ and $A_{b}$) on the SM-like Higgs boson mass, we impose $A_{t}=A_{b}$ with their magnitudes treated as free variables. Non-critical SUSY-breaking parameters are fixed: $m_{A}=2~{}{\rm TeV}$, $M_{1}=1~{}{\rm TeV}$, $M_{2}=2~{}{\rm TeV}$, $M_{3}=3~{}{\rm TeV}$. All parameters are defined at the renormalization scale $Q_{input}=1~{}{\rm TeV}$.
@@ 3.1 Research Strategy
In the specific parameter space scan, we set the value of $n_{\text{live}}=8000$. Through multiple trial-and-error processes, we choose the GNMSSM parameter space as shown in Table 1 . The likelihood function $\mathcal{L}$ that guided our scan is influenced by relevant experimental constraints. When the parameter points satisfy the experimental limits, $\mathcal{L}=1$; otherwise, $\mathcal{L}=\exp[100]$. The basic constraints considered include:
  • The SM-like Higgs boson: In our study, we consider $h_{1}$ or $h_{2}$ as the SM-like Higgs boson $h$. The mass of $h$ has an uncertainty of 3 GeV, so it should be restricted within the range 122 GeV $\leq m_{h}\leq$ 128 GeV.
  • Higgs data fit: Assuming that $h$ corresponds to the Higgs boson discovered by the LHC, its properties should align with the measurements reported by the ATLAS and CMS collaborations at the 95% confidence level. A p-value larger than 0.05 was required, which was tested by the code HiggsSignals-2.6.2 [64, 65, 66, 67].
  • Extra Higgs searches: Additional Higgs bosons must satisfy the limits from the direct searches at LEP, Tevatron, and LHC, which are evaluated using the code HiggsBounds-5.10.2 [68, 69, 70, 71, 72].
  • $B$-physics observables: Theoretical predictions for the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are required to be consistency with experimental measurements within $2\sigma$ [73].
Beyond the basic constraints, we incorporate constraints from HiggsTools [45], which integrates updated versions of HiggsBounds and HiggsSignals. These subpackages compare predictions against extra scalar searches (HiggsBounds) and 125 GeV Higgs measurements (HiggsSignals), respectively. Within HiggsTools, HiggsBounds thus provides direct constraints from extra scalar searches, while HiggsSignals assesses compatibility with the observed 125 GeV Higgs properties.
@@ 3.2 HiggsTools constraints
Figure 1: Surviving parameter samples after applying the basic constraints in Sec.3.1 are projected onto the plane of $\sigma(h)/\sigma_{SM}(h)Br(h\to a_{1}a_{1}\to bb\tau\tau)$ versus $m_{a_{1}}$. Blue points represent the surviving samples further constrained by HiggsTools (HT), while the black solid line denotes the ATLAS exclusion limit for the $h\to a_{1}a_{1}\to bb\tau\tau$ channel [38]. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively.
@@
In the light Higgs scenario, ATLAS and CMS direct search constraints, based on decay product observations, are largely implemented in HiggsTools. To compare experimental limitations on parameter space, we project surviving samples satisfying the basic constraints in Sec. 3.1 onto two-dimensional planes with vertical axes representing $\mu_{bb\tau\tau}\equiv\sigma(h)/\sigma_{\text{SM}}(h)\times\text{Br}(h\to a_{1}a_{1}\to bb\tau\tau)$, as shown in Figure 1. Left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively. The blue points represent the surviving samples further constrained by HiggsTools. We further incorporate ATLAS constraints [38] on $h\to a_{1}a_{1}\to bb\tau\tau$ (solid line), excluding parameter space above this limit. Comparative analysis reveals that the upgraded HiggsTools imposes significantly stronger constraints than legacy HiggsSignals/HiggsBounds, excluding a large portion of parameter regions. And its exclusion power further surpasses even the latest direct search limits (solid line), which rely on single-channel analyses. After applying the constraints from HiggsTools, the values of $\mu_{bb\tau\tau}$ can reach to 1.9% in the $h_{1}$ scenario and 2.1% in the $h_{2}$ scenario, respectively.
@@
Figure 2: Parameter samples excluded by HiggsTools in Figure 1 are color-coded based on the exclusion source. Red diamond points denote samples excluded solely by HiggsSignals (HS), orange square points represent those excluded exclusively by HiggsBounds (HB), and magenta pentagon points indicate samples excluded by both HB and HS.
@@
To compare the constraints from HiggsBounds and HiggsSignals on the GNMSSM parameter space in the light Higgs scenario, the parameter samples excluded by HiggsTools in Figure 1 are color-coded in Figure 2 based on the exclusion source. Red diamond points denote samples excluded solely by HiggsSignals, orange square points represent those excluded exclusively by HiggsBounds, and magenta pentagon points indicate samples excluded by both tools. Figure 2 demonstrates that HiggsBounds exhibits significantly stronger exclusion power than HiggsSignals. Parameter points excluded solely by HiggsSignals (red diamonds) are nearly absent, while points excluded by HiggsSignals are almost invariably also excluded by HiggsBounds, as indicated by the dominance of magenta points.
@@
In the $h_{1}$ scenario, parameter points with relatively large values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ (specifically, $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)\geq 2.5\%$) are excluded by both HiggsSignals and HiggsBounds. Points with smaller branching ratios are excluded only by HiggsBounds. This distinction arises because HiggsBounds imposes direct constraints, while HiggsSignals imposes indirect constraints on non-SM-like Higgs boson searches. Specifically speaking, an excessively large branching ratio $\text{Br}(h\to a_{1}a_{1})$, if sufficiently large, can affect the properties of the 125 GeV Higgs boson and this implies a large $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$, although the reverse is not necessarily true. However, small values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ always correspond to small $\text{Br}(h\to a_{1}a_{1})$, the latter having minimal impact on the properties of the 125 GeV Higgs boson.
@@
Figure 3: Similar to the right plot in Figure 2, here the label of the horizontal axis is $\text{Br}(h_{2}\to h_{1}h_{1}\to\tau\tau bb)$.
@@
Consistent with the $h_{1}$ scenario, analogous characteristics are also observed in the $h_{2}$ scenario, which simultaneously manifests distinctive features. The right panel of Figure 2 shows that some parameter points with $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$ (indicated by magenta points) are also excluded by HiggsSignals. This exclusion occurs because the exotic decay channel $h\to h_{1}h_{1}$ becomes kinematically accessible in this scenario, with an enhanced branching fraction. This enhancement may influence the properties of the SM-like Higgs boson and consequently affect the constraints imposed by HiggsSignals. Figure 3 plots the relationship between $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ and $Br(h\to h_{1}h_{1}\to\tau\tau bb)$ specifically in the $h_{2}$ scenario. It reveals an anticorrelation between these branching ratios across the parameter space, thereby validating the proposed explanation.
Figure 4: Similar to Figure 1, this plot exhibits the relationship between the parameter $\lambda$ and $\kappa$.
Figure 5: Like Fig. 1, this plot shows the SM-like Higgs boson component in the $h_{1}$ scenario.
Figure 6: Like Fig. 1, this figure shows the $h_{2}$ component (left panel) and the $h_{1}$ component (right panel) in the $h_{2}$ scenario.

### v3

@@ 2.4 Trilinear Higgs couplings in the GNMSSM
For the light Higgs scenario in this study, we assumed a light CP-odd Higgs boson ($a_{1}$) that satisfied $2m_{a_{1}}<m_{h}$ to enable the exotic decay $h\to a_{1}a_{1}$ (i.e. $h\to A_{s}A_{s}$) in both $h_{1}$ and $h_{2}$ scenarios. For the $h_{2}$ scenario, where the lightest CP-even Higgs boson $h_{1}$ is singlet-dominant, $h\to h_{1}h_{1}$ (i.e., $h\to h_{s}h_{s}$) opens when $2m_{h_{1}}<m_{h}$. Both ATLAS and CMS collaborations have searched for non-SM Higgs bosons $H$, $h_{s}$, $A_{H}$ and $A_{s}$, which established stringent exclusion bounds on key parameters, including the masses and relative couplings [61, 62].
The branching ratios of exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ are governed by their respective trilinear Higgs couplings $C_{ha_{1}a_{1}}$($\equiv C_{hA_{s}A_{s}}$) and $C_{hh_{1}h_{1}}$ ($\equiv C_{hh_{s}h_{s}}$). In the GNMSSM, these couplings proportionally scale to vacuum expectation values ($v_{u}$, $v_{d}$, $v_{s}$) or soft SUSY-breaking trilinear terms ($A_{\lambda}$, $A_{\kappa}$). Ref. [18] provides general expressions, but we explicitly define specific trilinear couplings that are relevant to our analysis below [35]:
@@
where the last terms in the preceding equations, which are suppressed by Higgs mixings, depend on $\lambda,\kappa,\tan\beta,v_{s},A_{\lambda}$ and $m_{N}$.
@@ 3 Numerical result
We constructed the GNMSSM model file using SARAH [63, 64, 65, 66] and generate particle spectra with SPheno [67, 68, 69, 63]. To explore the parameter space, we used the parallelized MultiNest algorithm [70], which uses nested sampling to identify high-likelihood regions, detect multiple solution modes, and compute Bayesian evidence.
@@
$\kappa$ Flat 0.75-0.750.750.75 $\tan\beta$ Flat 556060
$\lambda$ Flat 00.750.75 $v_{s}/{\rm TeV}$ Flat 0.10.11.01.0
$\delta$ Flat 1.0-1.01.01.0 $m_{N}/{\rm TeV}$ Flat 1.0-1.01.01.0
$m_{B}/{\rm GeV}$ Flat 11300300 $m_{C}/{\rm GeV}$ Flat 1.01.0300300
$A_{t}/{\rm TeV}$ Flat 1.01.03.03.0 $\mu_{\rm tot}/{\rm TeV}$ Flat 0.20.21.01.0
$A_{\kappa}/{\rm TeV}$ Flat 2.0-2.02.02.0
Table 1: Parameter space explored in this study. All input parameters used flat distributions based on their unambiguous physical interpretations. Considering the substantial radiative corrections that the third-generation squark trilinear couplings ($A_{t}$ and $A_{b}$) impose on the SM-like Higgs boson mass, we imposed $A_{t}=A_{b}$, where their magnitudes were treated as free variables. Non-critical SUSY-breaking parameters were fixed: $m_{A}=2~{\rm TeV}$, $M_{1}=1~{\rm TeV}$, $M_{2}=2~{\rm TeV}$, $M_{3}=3~{\rm TeV}$. All parameters were defined at the renormalization scale $Q_{input}=1~{\rm TeV}$.
@@ 3.1 Research strategy
In the specific parameter space scan, we set the value of $n_{\text{live}}=8000$. Through trial-and-error processes, we selected the GNMSSM parameter space as shown in Table 1. The likelihood function $\mathcal{L}$ that guided our scan is influenced by experimental constraints. When the parameter points satisfy the experimental limits, $\mathcal{L}=1$; otherwise, $\mathcal{L}=\exp[100]$. The basic constraints that we considered include:
  • The SM-like Higgs boson: In our study, we considered $h_{1}$ or $h_{2}$ as the SM-like Higgs boson $h$. The mass of $h$ has an uncertainty of 3 GeV, so it should be restricted in the range of 122 GeV $\leq m_{h}\leq$ 128 GeV.
  • Higgs data fit: Assuming that $h$ corresponds to the Higgs boson discovered by the LHC, its properties should align with the measurements reported by the ATLAS and CMS collaborations at the 95% confidence level. A p-value higher than 0.05 was required, which was tested by the code HiggsSignals-2.6.2 [71, 72, 73, 74].
  • Extra Higgs searches: Additional Higgs bosons must satisfy the limits from the direct searches at the LEP, Tevatron, and LHC, which were evaluated using the code HiggsBounds-5.10.2 [75, 76, 77, 78, 79].
  • $B$-physics observables: Theoretical predictions for the branching ratios of $B_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$ are required to be consistent with experimental measurements within $2\sigma$ [80].
Beyond the basic constraints, we incorporate constraints from HiggsTools [51], which integrates the updated versions of HiggsBounds and HiggsSignals. These subpackages compare predictions against extra scalar searches (HiggsBounds) and 125-GeV Higgs measurements (HiggsSignals), respectively. In HiggsTools, HiggsBounds provides direct constraints from extra scalar searches, whereas HiggsSignals assesses the compatibility with the observed 125-GeV Higgs properties.
@@ 3.2 HiggsTools constraints
Figure 1: Surviving parameter samples after we applied the basic constraints in Sec. 3.1 were projected onto the plane of $\sigma(h)/\sigma_{SM}(h)Br(h\to a_{1}a_{1}\to bb\tau\tau)$ versus $m_{a_{1}}$. The blue points represent the surviving samples after further constraints by HiggsTools (HT); the black solid line denotes the ATLAS exclusion limit for the $h\to a_{1}a_{1}\to bb\tau\tau$ channel [44]. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively.
@@
In the light Higgs scenario, the direct search constraints of ATLAS and CMS, which are based on decay product observations, were largely implemented in HiggsTools. To compare experimental limitations on the parameter space, we projected the surviving samples that satisfied the basic constraints in Sec. 3.1 onto two-dimensional planes, where the vertical axes represent $\mu_{bb\tau\tau}\equiv\sigma(h)/\sigma_{\text{SM}}(h)\times\text{Br}(h\to a_{1}a_{1}\to bb\tau\tau)$, as shown in Figure 1. The left and right panels correspond to the $h_{1}$ and $h_{2}$ scenarios, respectively. The blue points represent the surviving samples after further constraints of HiggsTools. We incorporated the ATLAS constraints [44] on $h\to a_{1}a_{1}\to bb\tau\tau$ (solid line) and excluded the parameter space above this limit. Comparative analysis reveals that the upgraded HiggsTools imposes stronger constraints than the legacy HiggsSignals/HiggsBounds. Approximately 96.7% (95.1%) of the parameter points were excluded in the $h_{1}$ ($h_{2}$) scenario. In addition, its exclusion power surpasses even the latest direct search limits (solid line), which rely on single-channel analyses. After the constraints from HiggsTools had been applied, the values of $\mu_{bb\tau\tau}$ could reach 1.9% in the $h_{1}$ scenario and 2.1% in the $h_{2}$ scenario.
@@
Figure 2: Parameter samples excluded by HiggsTools in Figure 1 are color-coded based on the exclusion source. The red diamond points denote samples that only HiggsSignals (HS) excludes, the orange square points denote samples that only HiggsBounds (HB) excludes, and the magenta pentagon points denote samples that both HB and HS exclude.
@@
To compare the constraints from HiggsBounds and HiggsSignals on the GNMSSM parameter space in the light Higgs scenario, the parameter samples excluded by HiggsTools in Figure 1 are color-coded in Figure 2 based on the exclusion source. The red diamond points denote samples that only HiggsSignals excluded, the orange square points denote those that only HiggsBounds excluded, and the magenta pentagon points indicate samples that both tools excluded. Figure 2 demonstrates that HiggsBounds exhibits stronger exclusion power than HiggsSignals. Almost no parameter points were only excluded by HiggsSignals (red diamonds), while points excluded by HiggsSignals are almost invariably also excluded by HiggsBounds, as indicated by the dominance of magenta points.
@@
In the $h_{1}$ scenario, parameter points with relatively large values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ (specifically, $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)\geq 2.5\%$) are excluded by both HiggsSignals and HiggsBounds. Points with smaller branching ratios are only excluded by HiggsBounds. This difference occurs because HiggsBounds imposes direct constraints, whereas HiggsSignals imposes indirect constraints on non-SM-like Higgs boson searches. Specifically, if an excessively large branching ratio $\text{Br}(h\to a_{1}a_{1})$ is sufficiently large, it can affect the properties of the 125-GeV Higgs boson and implies a large $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$, although the reverse is not necessarily true. However, small values of $\text{Br}(h\to a_{1}a_{1}\to\tau\tau bb)$ always correspond to small $\text{Br}(h\to a_{1}a_{1})$, the latter of which has minimal impact on the properties of the 125-GeV Higgs boson.
@@
Figure 3: Similar to the right plot in Figure 2, the label of the horizontal axis is $\text{Br}(h_{2}\to h_{1}h_{1}\to\tau\tau bb)$.
@@
Consistent with the $h_{1}$ scenario, analogous characteristics were observed in the $h_{2}$ scenario, which simultaneously manifested distinctive features. The right panel of Figure 2 shows that some parameter points with $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$ (indicated by magenta points) were also excluded by HiggsSignals, since the exotic decay channel $h\to h_{1}h_{1}$ becomes kinematically accessible in this scenario with an enhanced branching fraction. This enhancement may affect the properties of the SM-like Higgs boson and the constraints that HiggsSignals imposes. Figure 3 plots the relationship between $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ and $Br(h\to h_{1}h_{1}\to\tau\tau bb)$ in the $h_{2}$ scenario. It reveals an anti-correlation between these branching ratios across the parameter space, which validates the proposed explanation.
Figure 4: Relationship between parameters $\lambda$ and $\kappa$, similar to Figure 1.
Figure 5: SM-like Higgs boson component in the $h_{1}$ scenario, similar to Figure  1.
Figure 6: $h_{2}$ component (left panel) and $h_{1}$ component (right panel) in the $h_{2}$ scenario, similar to Figure  1.